Your friend, Suzie Whitson, has designed a new type of outdoor toy that helps children learn basic concepts such as colors, numbers, and shapes. Suzie’s product will target two groups: day care centers in warm climates and home school programs. Her company is Jiffy Jet and costs for last month follow: Factory rent $ 3,130 Company advertising 1,100 Wages paid to assembly workers 31,700 Depreciation for salespersons’ vehicles 2,150 Screws 570 Utilities for factory 885 Assembly supervisor’s salary 3,620 Sandpaper 180 President’s salary 5,110 Plastic tubing 4,160 Paint 270 Sales commissions 1,370 Factory insurance 1,070 Depreciation on cutting machines 2,130 Wages paid to painters 8,300 2. Determine the total cost for each of the following: a. Direct Materials b. Direct Labor c. Manufacturing Overhead d. Prime Cost e. Conversion Cost f. Total PRoduction Cost
Answer #1

a.

Direct material = Sandpaper + Plastic tubing + Paint

= 180+4,160+270

= $4,610

b.

Direct labor = Wages paid to assembly workers + Wages paid to painters

= 31,700+8,300

= $40,000

c.

Manufacturing Overhead = Factory rent + Screws + Utilities for factory + Assembly supervisor’s salary + Factory insurance +  Depreciation on cutting machines

= 3,130+570+885+3,620+1,070+2,130

= $11,405

d.

Prime cost = Direct material + Direct labor

= 4,610+40,000

= 44,610

e.

Conversion cost = Direct labor + Manufacturing overhead

= 40,000+11,405

= $51,405

f.

Total production cost = Direct material + Direct labor + Manufacturing overhead

= 4,610+40,000+11,405

= $56,015

Kindly comment if you need further assistance. Thanks‼!